Home‎ > ‎Reports‎ > ‎Crew100A‎ > ‎

day09


Crew 100A Log Book for February 20, 2011

Photos of the day





Commander's Check-In Report

Ivo Ferreira Reporting

Crew Physical Status: Good

Time Departed/Returned from EVA: EVA #6, #7

Brief Narrative of Field Mission Results: Today was an hectic day here at the MDRS.

Since yesterday the rain disrupted our schedule, we had to work harder to compensate it.

We started quite early with some of us conducting the usual habitability experiments while the others checked the terrain conditions and the site objectives for the SALM Mission (Site Analof Lune Mars Group) involving a Drone Flyby, iteration with science group in La Reunion to redefine the objectives, human EVA and shipping of

samples.

will work on stabilizing the high gain antenna on top of the car next Before breakfast we still hurried to test the Wifi relay station that we developed but this time using a car as the mobile unit providing the necessary power (no more batteries). We managed to drive Max (Rover) almost until the objective (Dino Site about 700m from the Habitat) but the connection was disrupted and the mission aborted. We time.

After lunch it was time to finish the Tiger experiment with the final human EVA (EVA#6) to pick up samples (of both terrain and rocks) at the two outcrops of white and red rocks, previously visited by the drone and the rover on Feb 18th. The mission was a complete success and is now in the final stage of analyzing physically and chemically the samples.

Then it was time to another EVA to test the effect of stress in the communication (EVA#7) involving three of the crew members (one at the mission control and two other on the field), this time with more complex mission objectives (locomotion by ATV).

Finally, at the end of the day the wind calmed down and we finally managed to conduct the drone flybys on the two sites for the SALM Mission.

Engineering/Hab Maintenance: Problem with leak of the hot water system fixed.

Plans for Tomorrow: Proceed with SALM mission. Develop stabilizing system for Max WiFi relay station. Continue with experiments: habitability, and EVAs to test stress level in the afternoon.

Reports scheduled for today: Engineering, Commander, Journalist, Science, EVA #6 and EVA#7

Inventory: None

Support Requested: N/A

Miscellaneous: N/A


Commander's Report

Ivo Ferreira Reporting

After the rain, a hectic day

What a day... Since yesterday we had a "forced" relaxing day, today we had to shift  gears into our fastest speed to get to the plan. I know it is Sunday back on Earth but here in Mars the week has 7 working days, double shifts with no holidays.

We started our day quite early to continue with Ayako's experiments and check out the terrain for one of the complex missions we have to gather with the SALM group in La Reunion. Romain and I left to search the sites but came back with more mud than pictures.

The Martian GPS system is still under development, so we get errors of up to 40 meters in the GPS here at the station. We had to do it "old style" with visual identification of the targets which is an extremely difficult task when there are now three small rivers that did not exist until yesterday.

Before lunch there was still time to test the range of Max (the rover) to get to a Dino site nearby, it was a true effort of coordination between me, Kévin and Audrey. This time we decided to not rely on the old junkyard batteries and use our portable vehicle (AKA Ford Expedition) to place (and power) the relay station.

Lunch time and it already felt like a full day of work...  but it was time for the second shift.

Ayako and I started the afternoon by finishing the Tiger mission (from 18th of February) by picking up several samples at two rock outcrops nearby. These have been confirmed by a previous parrot flyby and a rover teleoperation (that ended under "artificial light"). It is

now time for Audrey to analyze the samples with some optical equipment and low pH buffer solution (as a substitute for HCl).

Romain was already suited up when we came back and departed on a mission on the ATV with Ayako. After teaching Ayako a  5-minute course on driving in the desert full of mud at high speeds, they went on to save the planet on a really stressful (simulated) mission.

It was already almost 6pm but wind had stopped and we could not resist. It was time to use the Parrot and accomplish the mission we had started the day planning.

Uff... Time for dinner... (luckily it is a cooking day).


Engineering Report

Kévin Theatre Reporting

Generator/Electricity:

· Kitty (Diesel Generator) (Used? Yes/No): Yes

· Honey (Gas Generator) (Used? Yes/No): No

· Battery Charge Level: Checked by DG

o 10.59 V everstart

o 12.37 V armor plate

o 13.32 V green


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is good.


Fuel Status:

· Diesel (Full/0.75/0.5/0.25/Empty): 0.76

· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.54

· Gasoline Tank (Full/0.75/0.5/0.25/Empty): 11 Gallons


Notes/Comments:


Oil:

· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

· Honey Oil Quantity (# of Quarts in storage): Leave it to DG

· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

· Car Oil Quantity (# of Quarts in storage):


Water Status:

· Outside Potable Water Tank Level (inches from bottom): 17

· Trailer Potable Water Tank Level (inches from bottom): Full

· Hab Potable Water Tank Level (inches from bottom): 11

· Potable Water Meter Reading: 423451

Notes/Comments:


GreenHab:

· Condition of plants in Tank 1: green water

· Condition of plants in Tank 2: green things

· Condition of plants in Tank 3: green things

· Greenhab notes/ comments: The level of water of the 3 tanks has reduced a little.


Observatory:

· Telescope Used? (Yes/no) No

Observatory Notes:


Transportation:

Hab Car (NOT RENTAL)

· Used (yes/no): yes

· Oil Checked (yes/no): No

· Tire Status: checked

· Odometer at end of day: 123614

· Notes/Comments on Hab Car: just to supply electricity to relay for the Rover during our first attempt to do mission SALM.


ATV 1 (Opportunity):

· Used: Yes

· Oil Checked: No

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


ATV 2 (Spirit):

· Used: No

· Oil Checked: No

· Fuel Consumed:

· Tire Status: filled

· Comments: Still not fixed


ATV 3 (Viking I):

· Used: No

· Oil Checked: No

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


ATV 4 (Viking II):

· Used: yes

· Oil Checked: no

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


Heating and Ventilation:

· Thermostat setting upstairs (Farenheit): 70

· Actual temperature upstairs (Farenheit): 72

· Thermostat setting downstairs (Farenheit): 65

· Actual temperature downstairs (Farenheit): 63


Computers/Networking Infrastructure:

Notes/comments on internet/computers?:


General Engineering

1. Notes/Comments:

DG fixed our hot water leak.


EVA 6 Report

Ivo Ferreira Reporting

EVA #6, 15:30 - 16:10

crew members : Ivo Ferreira, Ayako Ono

Site location : 38.40617, -110.791054

Transit mode : Foot

Objective : Finish Tiger Mission with collection of samples at two rock outcrops (White and Red Rocks)

Result : We managed to gather 2 samples for each of the outcrops (one rock and a soil sample for each outcrop).

Lesson learned : During an EVA it is important to define a strategy if the communications go down (hand signs). The choice of the tools is quite important. One of the persons should collect the sample while the other provides the tools to avoid contamination.


EVA 7 Report

Romain Benchenafi Reporting

EVA #7, 16:54 - 17:10

Crew members : Ivo FERREIRA, Ayako ONO, Romain BENCHENAFI

Site location :  UTM 12S 0519141 4251993

Transit mode : ATV

Objective : experiment on communication related to stress

Result : Managed to comply with the objective in spite of the constraints.

Lesson learned : It is difficult to follow the GPS and communicate without the proper equipment while driving an ATV.


Journalist Report

Romain Benchenafi Reporting

It is the beginning of our second week of simulation and fortunately for us we have been able to continue our experiments with the return of the sun. Only the wind still interferes in our schedule.

We also have to highlight the success of mission TIGER and the start of an investigation for the SALM institute.


Chef's Report

Ivo Ferreira Reporting

1a) Today was a non-cooking day

1b) Was today a special day celebrated at the main meal with special food or activity?

No

2)  List the foods served at today's main meal (usually the evening meal), giving full names of each. Fried Tofu, Potatoes au Gratin, Oreganos, Bacon

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

  Nothing

4)  Recipes for anything more complicated than following package directions exactly, or rehydrating.  Everything followed the instructions on the package.

5) Lessons learned  N/A

6) Comments/questions for Kim and Jean, the food study investigators

What is the different between non-cooking and pre-prepared food?